A fan shuts down, and its angular speed decreases from 10.0 rad / s to 6.3 rad / s in 5.0 s. What is the magnitude of the angular acceleration of a fan?
Answer #1

Angular acceleration:

A = (W1 - W2) / t = (6.3-10) /5

= -0.37 rad/s2

Note : negative sign is for deceleration